OpenStudy (anonymous):

Help please, derivative of 5csc[sin^2(x^2+1)]

OpenStudy (anonymous):

sorry sorry, it must be sin^2(x^2+1)

OpenStudy (dls):

Let us say \[\LARGE y=5\csc[\sin^2(x^2+1)]\] We therefore need to compute dy/dx.U Use the chain rule :) \[ \frac{dy}{dx}=-5 \sin(\sin^2(x^2+1)) \times 2\sin(x^2+1) \times \cos(x^2+1) \times 2x \]

OpenStudy (anonymous):

can you teach me please? step-by-step? if it's ok...